OpenStudy (mayliz):

f(x)=3x^2+4x+1 f(x)=3(x+a)^2+b solve a and b

OpenStudy (mayliz):

i have got a=2/3 b=-1/3 Is this right ?

OpenStudy (nincompoop):

are they both f(x)?

OpenStudy (mayliz):

yeah 3(x+a)^2+b=3x^2+4x+1

OpenStudy (nincompoop):

expand 3(x+a)^2 and 3x^2 first

OpenStudy (nincompoop):

I guess I want to see how you came up with your answers.

OpenStudy (mayliz):

3x^2+6xa+3a^2+b=3x^2+4+1 6a=4 ............ a=2/3 3a^2+b=1 .....b=-1/3

OpenStudy (nincompoop):

2/3 and -1/3 it is

OpenStudy (anonymous):

both function are same .so we can write as bellow 3(x+a)^2=3x^2+6xa+3a^2 so 6xa=4x a=2/3 and 3a^2+b=1 3*(2/3)^2+b=1 b=-1/3